In a 10-item exam, of which only 6 items are to be solved, how many ways can a student answer the exam?
Answer by Hawksfan(61)

Use the formula
P = n!/(n- r)!

Example 5! = 5*4*3*2*1 for factorial !


= 10! / (10-6)!

10! / 4!
= (10*9*8*7*6*5*4!)/ 4!

= 90*56*30 = 151,200 ways